shila is 15 years older than her brother sanjay. ten years from now shila will be twice as sanjay. find their present ages.
Answer by Maths68(1474)

Let
Present age of Shila = h
Present age of Sanjay = s
Shila is 15 years older than Sanjay
h=s+15............(1)

Ten years after
Shila will be = h+10
Sanjay will be = s+10
Then
Shila will be twice as Sanjay
(h+10)=2(s+10)
h+10=2s+20
h=2s+20-10
h=2s+10...................(2)
Substitute the value of h from (1) to (2)
s+15=2s+10
s-2s=10-15
-s=-5
Cancel out (-)
s=5
Put the value of s in (1)
h=s+15...........(1)
h=5+15
h=20


Present age of Shila = h = 20
Present age of Sanjay = s = 5